# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v1

@@ 6.1. GUI Classification
GUI classification aims to categorize a screenshot as a whole under a specific label. It can be the fundamental task for other GUI-related tasks, like GUI captioning (Leiva et al., 2023), or the association of user screenshots (e.g. submitted in support requests) to certain system requirements and components (Maalej et al., 2009; Roehm et al., 2013; Haering et al., 2021). In this section, we explore the performance of UIClip on GUI classification under zero-shot and linear probe setting using the Enrico dataset (Leiva et al., 2020).
Figure 9. Overview of zero-shot and linear probe classification on Enrico dataset with UIClip (left side is adapted from (Radford et al., 2021))
@@
Overview of zero-shot and linear probe classification on Enrico dataset with UIClip (left side is adapted from (Radford et al., 2021))
@@ 6.1.1. Model Architecture
As performed by Radford et al. (Radford et al., 2021), we evaluated the performance of UIClip and baselines on GUI classification under zero-shot setting and linear probe setting as shown on Figure 9. Under both settings, the image encoder is frozen.
  • Zero-shot. In the zero-shot process, all classification labels are embedded to text embedding with the text encoder. During the classification process, the screenshot image is encoded by the image encoder, and the resulting screenshot embedding is compared to all text label embeddings. The label exhibiting the highest similarity to the GUI image embedding is subsequently declared as the predicted outcome.
  • Linear Probe. While in the linear probe setting, we added a linear layer of shape $m\times k$ to the output of UIClip image encoder. m refers to the output dimension of image encoder, while k is the number of classification labels.
@@ 6.1.2. Dataset
We use the Enrico dataset to evaluate the performance of GUI classification. Leiva et al. (Leiva et al., 2020) revised a random sample of 10k screenshots from Rico dataset to create Enrico (shorthand of Enhanced Rico): a human annotated dataset comprising 1460 screenshots and 20 design topics, including camera, chat, media player, etc.
In the linear probe setting, the Enrico dataset was divided into an 80:20 ratio to establish a training set and a test set, which were used respectively for the model training and testing. For the zero-shot setting, all available data was employed solely for testing purposes.
6.1.3. Baselines
We compared our model in GUI classification with the following baselines, performing both zero-shot classification and linear probe classification:
  • CLIP. We also evaluated the performance of CLIP on GUI classification. The setting of zero-shot and linear probe evaluation of CLIP is identical to the setting of UIClip. Their only difference is the parameters’ weights of the two models.
6.1.4. Training Details of Linear Probe
For the training and evaluation of linear probe classifier, we split the dataset into training and test set in a ratio of 80:20 in a stratified manner. During training, the image encoder is frozen, the output embedding is served as input of the final linear layer. We trained the model for 100 epochs using AdamW optimizer with an initial learning rate $1e^{-5}$, batch size 128. We conducted a ten-fold cross-validation, which involved randomly splitting the training and testing sets ten instances, and subsequently computing the average performance spanning across these runs.
@@ 6.1.5. Evaluation Metric
The evaluation of the classifiers is performed by utilizing precision, recall, and the F1 score. Given the relatively small size of the Enrico dataset, especially when compared to the 20 available classification labels, we opt not to calculate the precision, recall, and F1 score for each individual label. In contrast, we calculate their weighted average.
@@
Zero-shot Linear Probe
@@ 6.1.6. Results and Discussion
As shown in Table 6, vision-language models, specifically CLIP and UIClip, clearly outperform OCR+BGE as expected. This reaffirms the assertion that mere analysis of text displayed on screenshots is insufficient for substantial GUI understanding, making the vision information crucial. Notably, UIClip achieves superior results compared to CLIP, and records an F1 score that exceeds that of CLIP by 20 percent in both settings. This highlights the advantage of our UIClip model for GUI classification tasks over the baseline models, given its enhanced knowledge about the GUI domain accumulated during the pre-training phase. Particularly, the Linear Probe led to encouraging precision and recall of over 60% (given the large number of classes in the classification task and the stright application of the model).
The efficacy of vision-language models, as measured under the linear probe setting, surpasses that determined under the zero-shot setting. This superior performance can be rationalized in terms of the additional knowledge imparted during the training phase about the Enrico dataset. However, a contrary pattern was found concerning OCR+BGE, which showed diminished performance following linear probe training. We hypothesize that this outcome is attributable to the text-only approach’s inadequate competency in extracting visual information from the screenshots, which renders the the training stage ineffective.
@@ 6.2. Sketch-to-GUI Retrieval
Rather than searching GUI by textual query, it is also possible to perform the retrieval using a UI sketch. Based on UIClip, we built a Dual UIClip ViT model for sketch-to-GUI retrieval. An evaluation on the Swire dataset (Huang et al., 2019) shows that our model outperform Swire and CLIP on sketch-to-GUI retrieval, with a recall@10 of 0.685.
Figure 10. Overview of the training and inference of Dual UIClip ViT model for sketch-to-GUI retrieval
@@
Overview of the training and inference of Dual UIClip ViT model for sketch-to-GUI retrieval
@@ 6.2.1. Model Architecture
In order to adapt UIClip for sketch-to-GUI retrieval, we employed its image encoder. The left side of the Figure 10 illustrates the architecture of the model, which encompasses a screenshot encoder and a sketch encoder. Both of these components are initialized from UIClip’s image encoder. During the training phase, the weights of the screenshot encoder are held constant, whilst training is solely conducted on the sketch encoder. While during the inference phase, which is illustrated at the right side of the Figure 10, both of the two encoders are frozen. The embedding of the query sketch is compared with the embeddings of each screenshot, the top-k screenshots whose embedding is the most similar with the sketch embedding are retrieved.
@@ 6.2.2. Dataset
The dataset employed in this study was procured from Swire (Huang et al., 2019), which encompasses 3802 sketches corresponding to 2201 screenshots from 167 apps within the Rico dataset. This dataset was compiled by four designers, who contributed to sketching 505, 1017, 1272, and 1008 screenshots respectively. However, it was discovered that certain images from this dataset were not present within the Rico dataset, hence after appropriate filtration, a refined dataset featuring a total of 3551 sketches was obtained. The split between the training and test sets within Swire’s dataset was not explicitly delineated in their work. Consequently, for the purposes of our study, we adopted 486 sketches from one designer as the test set, while the sketches from three other designers constituted the training set. This is to ensure that the model exhibits generalization across sketches developed by various designers.
@@ 6.2.3. Baselines
To investigate the benefit of using UIClip in sketch-to-GUI retrieval, we compared our model with the following baselines.
  • Swire. Swire establishes itself as the pioneer sketch-to-GUI retrieval model (Huang et al., 2019). This model comprises two distinct VGG-A networks (Simonyan and Zisserman, 2015) that independently compute the embeddings for screenshots and sketches. The retrieval process is subsequently effectuated by measuring the similarity between these embeddings. Given that Swire’s implementation is not publicly accessible, we took it upon ourselves to recreate it according to the details laid out in the original paper.
  • Dual CLIP ViT. The architecture of the Dual CLIP ViT is identical to that of UIClip, with their main distinguishing factor being their respective parameters’ weights. The weights of the image encoder of Dual CLIP ViT are derived directly from the CLIP model.
@@ 6.2.4. Training Details of Fine-tuning
Under zero-shot setting, the Dual UIClip ViT and Dual CLIP ViT have not been trained. While under fine-tuning setting, the Dual UIClip ViT and Dual CLIP ViT model has been trained in contrastive learning manner on the training set for 10 epochs. We used AdamW optimizer with an initial learning rate $1e^{-5}$, batch size 256.
For the training of Swire, we tried to use the same configurations as specified in the original paper, including a learning rate of 0.01 and a batch size of 32. However, we noticed a significant shortfall in performance with these settings, wherein the top-10 recall stood at a mere 0.02. To address this, we made the decision to utilize a batch size of 64 and a learning rate of $1e^{-5}$. The original article does not specify the number of training epochs employed; thus, we proceeded to train it for 100 epochs utilising the training set, until we observe convergence in the training loss.
@@ 6.2.5. Evaluation Metric
Similar as the evaluation of text-to-GUI retrieval presented on the Section 4.1.4, we use the recall@1, recall@3, recall@5 and recall@10 for evaluation.
@@ 6.2.6. Results and Discussion
Table 7 presents the evaluation outcomes of UIClip and various baseline models on the Swire dataset within both zero-shot and fine-tuning settings. Our observation revealed that even under zero-shot setting, the Dual UIClip ViT has a recall@10 score of 0.224, which indicates its valuable application in sketch-to-GUI retrieval. In contrast, the Dual CLIP ViT was not as effective, as evidenced by its low recall@10 score of 0.053.
Fine-tuning can significantly improve the models’ performance. After fine-tuning, the recall@10 score of the Dual CLIP ViT increased to 0.553, while the Dual UIClip ViT exhibited a great improvement, reaching the score of 0.772. These results substantiate the superior efficacy of UIClip in comparison to CLIP for sketch-to-GUI retrieval.
The performance of our implementation of Swire model, with a recall@1 of 0.027 and a recall@10 of 0.214, is not congruent with the metrics reported in the originating study by Huang et al. (Huang et al., 2019). The cited research demonstrated significantly higher recall values of 0.159 and 0.609, at ranks 1 and 10 respectively. This discrepancy could potentially be attributed to two factors: 1) Our evaluation used 486 pairings for testing, in contrast to the 276 pairings used in the original research. The size of the test set can dramatically influence retrieval performance outcomes; 2) The divergence in implementation methods could also account for differing results. It’s important to note, however, that the original Swire model’s implementation is not open-source for replication. Nonetheless, even if we take into consideration the results from Huang et al. (Huang et al., 2019), our Dual UIClip ViT model still demonstrates superior performance, with a score of 0.368 for recall@1 and 0.772 for recall@10.
6.3. Additional Tasks and Applications
Besides GUI classification and sketch-to-GUI retrieval, our model has the potential to be adapted for other GUI-related tasks.
6.3.1. GUI Captioning
GUI captioning is a fundamental task in GUI understanding, where the model predicts a textual informative caption to a given GUI (Wang et al., 2021; Li and Li, 2022; Leiva et al., 2023). ClipCap (Mokady et al., 2021) is an image captioning model. It applies CLIP as image encoder, followed by a simple mapping network to generate the caption prefix. Subsequently, this caption prefix is utilized for the production of the caption through a language model. Given the superior performance of the UIClip model in understanding GUI information, it is possible to enhance ClipCap on GUI Captioning task by replacing the CLIP image encoder with the image encoder of UIClip.
6.3.2. GUI generation
GUI generation is the task that generates GUIs that aligns to the given specifications. Wei et al. (Wei et al., 2023) introduced UI-Diffuser that leverage stable diffusion model to generate mobile GUIs through simple textual descriptions and UI components. The text encoder of CLIP is the core component of the stable diffusion model, it converts the textual description to embedding that can be processed by the subsequent networks. However, given that the CLIP model is trained on general image-caption pairs, it may fall short on understanding the GUI-related descriptions. A potential improvement to the stable diffusion model for GUI generation tasks could be the replacement of the text encoder used in CLIP with that of UIClip’s.
@@ 7.1. Mobile GUI Retrieval
GUI retrieval refers to the search of existing GUI designs with queries of various format.
By wireframe: Wireframe is an image that represents the skeletal layout of a screenshot. Given a wireframe image, one can retrieve the corresponding screenshots. Deka et al. (Deka et al., 2017) and Liu et al. (Liu et al., 2018) trained autoencoder (Bengio, 2009) by reconstructing the wireframe to generate the wireframe embedding for GUI retrieval by embedding similarity. Chen et al. (Chen et al., 2020) performed wireframe-based GUI retrieval by training an autoencoder for encoding the visual semantics of UI designs using a large database of UI design wireframes.
@@
By sketch image: UI sketch is often used at the early stage of design, it can also be used as the query for the GUI retrieval. Huang et al. introduced Swire (Huang et al., 2019), which comprises two separate VGG-A networks (Simonyan and Zisserman, 2015). These networks calculate the embeddings for screenshots and sketches, respectively. The retrieval process is subsequently performed by measuring the similarity of these embeddings. Unlike Swire that treats the sketch image as a unit, Mohian et al. (Mohian and Csallner, 2022, 2023) performed object detection on sketch to identify the UI components on a sketch image, and then find the screenshots in Rico dataset whose type and location of UI components match best the sketch.
@@
By screenshot image: Using a screenshot from an app, it is possible to retrieve similar screenshots from the GUI repository. Screen2vec (Li et al., 2021) extracts the UI components, layouts and app description to generate a screen embedding, querying similar screenshots is done by comparing the similarity of screen embedding. On the other hand, VINS (Bunian et al., 2021) creates screen embedding by combining both image embedding and UI components embedding for screen searching. Instead of querying screenshots, GUIFetch (Behrang et al., 2018) searches for apps in public repositories that are as similar as possible to the provided app’s screens and transitions between them.
@@
By textual query: There are also other works that use textual queries for GUI retrieval. GUIGLE (Bernal-Cardenas et al., 2019) has indexed a collection of GUIs along with their metadata from the ReDraw dataset (Moran et al., 2020). GUIGLE leverages various information sources such as text displayed on the screen, name of UI components, and app name to retrieve relevant screens based on the keyword matching to the given query. On the other hand, RaWi (Kolthoff et al., 2023) uses a BERT-based ranking model to retrieve GUIs from the Rico dataset (Deka et al., 2017). This retrieval process involves matching the GUI text (text displayed on the screen, GUI activity name and GUI component identifier) with the textual query in order to identify the most relevant images. Unlike the preceding studies that focus on retrieving entire screenshot, Gallery D.C. (Chen et al., 2019b; Feng et al., 2022) allow users to search UI components, like button and checkbox. It provide a GUI gallery that allow users to search UI components with different filters, including component type, size, color, app category and displayed text. The aforementioned methods used the textual query to search the text presented on the GUIs, while the non-textual information have not been taken into account.

### v2

@@ 6.1. GUI Classification
GUI classification aims to categorize a screenshot as a whole under a specific label. It can be a fundamental activity for other GUI-related tasks, like GUI captioning (Leiva et al., 2023), the tagging of user screenshots (e.g. submitted in support requests) with certain requirements, components, or devices (Maalej et al., 2009; Roehm et al., 2013; Martens and Maalej, 2019a), as well as an early prototyping by a larger group and multiple screens (Pham et al., 2018). In this section, we explore the performance of GUIClip on GUI classification under zero-shot and linear-probe setting using the Enrico dataset (Leiva et al., 2020).
Figure 9. Overview of zero-shot and linear-probe classification on Enrico dataset with GUIClip (left side is adapted from (Radford et al., 2021)).
@@
Overview of zero-shot and linear-probe classification on Enrico dataset with GUIClip (left side is adapted from (Radford et al., 2021)).
@@ 6.1.1. Model Architecture
Like Radford et al. (Radford et al., 2021), we evaluated the performance of GUIClip for GUI classification under zero-shot and linear-probe settings as shown on Figure 9. Under both settings, the image encoder is frozen.
  • Zero-shot. In the zero-shot setting, all classification labels are embedded to text embedding with the text encoder. During the classification process, the screenshot image is encoded by the image encoder, and the resulting screenshot embedding is compared to all text label embeddings. The label exhibiting the highest similarity to the GUI image embedding is subsequently declared as the predicted outcome.
  • Linear-probe. In the linear-probe setting, we added a linear layer of shape $m\times k$ to the output of GUIClip image encoder. m refers to the output dimension of image encoder, while k is the number of classification labels.
6.1.2. Baselines
We compared GUIClip performance with two baselines for zero-shot and linear-probe classification:
  • CLIP. We also evaluated the performance of CLIP on GUI classification. The setting of zero-shot and linear-probe evaluation of CLIP is identical to the setting of GUIClip. Their only difference is the parameters’ weights of the two models.
@@ 6.1.3. Dataset
We used the Enrico dataset to evaluate the model’s performance for GUI classification. Leiva et al. (Leiva et al., 2020) used a random sample of 10k screenshots from Rico dataset to create Enrico (shorthand of Enhanced Rico): a human annotated dataset comprising 1460 screenshots and 20 design topics, including camera, chat, media player, etc. In the linear-probe setting, the Enrico dataset was divided into an 80:20 ratio to establish a training set and a test set. For the zero-shot setting, all available data was employed solely for testing.
6.1.4. Training Details of Linear-probe Classification
For the training and evaluation of linear-probe classifier, we split the dataset into training and test set in a ratio of 80:20 in a stratified manner. During training, the image encoder is frozen, the output embedding is served as input of the final linear layer. We trained the model for 100 epochs using AdamW optimiser with an initial learning rate $1e^{-5}$, batch size 128. We conducted a 10-fold cross-validation, which involved randomly splitting the training and testing sets ten times, and subsequently computing the average performance spanning across these runs.
@@ 6.1.5. Evaluation Metric
The evaluation of the classifiers is performed based on precision, recall, and the F1 score. Given the relatively small size of the Enrico dataset, especially when compared to the 20 available classification labels, we did not calculate the precision, recall, and F1 score for each individual label. In contrast, we calculate their weighted average.
@@
Zero-shot Linear-probe
@@ 6.1.6. Results and Discussion
As shown in Table 6, vision-language models, specifically CLIP and GUIClip, clearly outperform OCR+BGE as expected. This reaffirms the assertion that mere analysis of text displayed on screenshots is insufficient for substantial GUI understanding, making the vision information crucial. Notably, GUIClip achieves superior results compared to CLIP, and records an F1 score that exceeds that of CLIP by 20 percent in both settings. This highlights the advantage of our GUIClip model for GUI classification tasks over the baseline models, given its enhanced knowledge about the GUI domain accumulated during the pre-training phase. Particularly, the Linear-probe led to encouraging precision and recall of over 60% (given the large number of classes in the classification task and the straight application of the model).
The efficacy of vision-language models, as measured under the linear-probe setting, surpasses that determined under the zero-shot setting. This superior performance can be explained by the additional knowledge gained during the training phase from the Enrico dataset. However, a contrary pattern was found concerning OCR+BGE, which showed diminished performance following linear-probe training. We hypothesise that this is because the text-only approach is inadequate to extract visual information from the screenshots, which makes the training stage ineffective.
@@ 6.2. Sketch-to-GUI Retrieval
Instead of searching for a textual query, it is also possible to perform the GUI search using a UI sketch. Based on GUIClip, we built a Dual GUIClip ViT model for sketch-to-GUI retrieval. An evaluation on the Swire dataset (Huang et al., 2019) shows that our model outperform Swire and CLIP on sketch-to-GUI retrieval, with a recall@10 of 0.685.
Figure 10. Overview of the training and inference of Dual GUIClip ViT model for sketch-to-GUI retrieval.
@@
Overview of the training and inference of Dual GUIClip ViT model for sketch-to-GUI retrieval.
@@ 6.2.1. Model Architecture
In order to adapt GUIClip for sketch-to-GUI retrieval, we employed its image encoder. The left side of Figure 10 illustrates the architecture of the model, which encompasses a screenshot encoder and a sketch encoder. Both of these components are initialised with the image encoder of GUIClip. During the training phase, the weights of the screenshot encoder are held constant, whilst training is solely conducted on the sketch encoder. During the inference phase, which is illustrated at the right side of the Figure 10, both encoders are frozen. The embedding of the query sketch is compared with the embeddings of each screenshot. The top-k screenshots whose embedding is the most similar with the sketch embedding are retrieved.
@@ 6.2.2. Dataset
The dataset employed in this study was procured from Swire (Huang et al., 2019), which encompasses 3802 sketches corresponding to 2201 screenshots from 167 apps within the Rico dataset. This dataset was compiled by four designers, who contributed to sketching 505, 1017, 1272, and 1008 screenshots respectively. However, it was discovered that certain images from this dataset were not present within the Rico dataset. Hence after appropriate cleaning, a refined dataset featuring a total of 3551 sketches was obtained. The split between the training and test sets within Swire was not explicitly delineated in their work. Consequently, for the purposes of our study, we adopted 486 sketches from one designer as the test set, while the sketches from three other designers constituted the training set. This ensured that the model learns to generalise across sketches developed by various designers.
@@ 6.2.3. Baselines
We compared GUIClip’s performance in sketch-to-GUI retrieval with the following baselines:
  • Swire. Swire establishes itself as the pioneer sketch-to-GUI retrieval model (Huang et al., 2019). This model comprises two distinct VGG-A networks (Simonyan and Zisserman, 2015) that independently compute the embeddings for screenshots and sketches. The retrieval process is subsequently conducted by measuring the similarity between these embeddings. Given that Swire’s implementation is not publicly accessible, and since we were unable to get it from the authors, we took it upon ourselves to recreate it according to the details in the original paper.
  • Dual CLIP ViT. The architecture of the Dual CLIP ViT is identical to that of GUIClip, except for their respective parameters’ weights. The weights of the image encoder of Dual CLIP ViT are derived directly from the CLIP model.
@@ 6.2.4. Training Details of Fine-tuning
Under zero-shot setting, the Dual GUIClip ViT and Dual CLIP ViT have not been trained. While under fine-tuning setting, the Dual GUIClip ViT and Dual CLIP ViT models have been trained with contrastive learning on the training set for 10 epochs. We used AdamW optimiser with an initial learning rate $1e^{-5}$, batch size 256.
For the training of Swire, we tried to use the same configurations as specified in the original paper, including a learning rate of 0.01 and a batch size of 32. However, we noticed a significant shortfall in performance with these settings, wherein the top-10 recall stood at a mere 0.02. To address this, we made the decision to utilise a batch size of 64 and a learning rate of $1e^{-5}$. The original article does not specify the number of training epochs employed. Thus, we proceeded to train it for 100 epochs utilising the training set, until we observe convergence in the training loss.
@@ 6.2.5. Evaluation Metric
Similar to the evaluation of text-to-GUI retrieval presented on the Section 4.1.4, we use recall@1, recall@3, recall@5 and recall@10.
@@ 6.2.6. Results and Discussion
Table 7 presents the evaluation outcomes for GUIClip and the baseline models on the Swire dataset within both zero-shot and fine-tuning settings. Our observation revealed that even under zero-shot setting, the Dual GUIClip ViT has a recall@10 score of 0.224, which indicates its potential for sketch-to-GUI retrieval. In contrast, the Dual CLIP ViT was not as effective, as evidenced by its low recall@10 score of 0.053.
Fine-tuning can significantly improve the models’ performance. After fine-tuning, the recall@10 score of the Dual CLIP ViT increased to 0.553, while the Dual GUIClip ViT exhibited a remarkable improvement, reaching 0.772. These results substantiate the superior efficacy of GUIClip in comparison to CLIP for sketch-to-GUI retrieval.
The performance of our Swire implementation, with a recall@1 of 0.027 and a recall@10 of 0.214, is not congruent with the metrics reported in the originating study by Huang et al. (Huang et al., 2019). The researchers reported a significantly higher recall values of 0.159 and 0.609, at ranks 1 and 10 respectively. This discrepancy could potentially be attributed to two factors: 1) Our evaluation used 486 pairings for testing, in contrast to the 276 pairings used in the original research. The size of the test set can dramatically influence retrieval performance outcomes; 2) The implementation details could also lead to differing results. It is important to note, however, that the original Swire model’s implementation is not available for replication. Nonetheless, even if we take into consideration the results from Huang et al. (Huang et al., 2019), our Dual GUIClip ViT model still demonstrates superior performance, with a score of 0.368 for recall@1 and 0.772 for recall@10.
@@ 7.1. Mobile GUI Retrieval
GUI retrieval refers to the search of existing GUI designs (usually images) with queries of various format.
@@
By sketch image: UI sketches are often used at the early stage of the design. They can also be used as a query. As detailed in Section 6.2, Huang et al. introduced Swire (Huang et al., 2019), which comprises two separate VGG-A networks (Simonyan and Zisserman, 2015) to compute the embeddings for screenshots and sketches, respectively. The retrieval process is subsequently performed by measuring the similarity of these embeddings. Unlike Swire that treats a sketch image as a unit, Mohian et al. (Mohian and Csallner, 2022, 2023) performed object detection on the sketches to identify their UI components, which are then used to find the screenshots in Rico dataset with the best matching types and locations of the UI components.
By wireframe: A wireframe is an image that represents the skeletal layout of a screenshot. Given a wireframe image, one can retrieve the corresponding screenshots. Deka et al. (Deka et al., 2017) and Liu et al. (Liu et al., 2018) trained an autoencoder (Bengio, 2009) for UI layout similarity, which supports query-by-example search over UIs. Chen et al. (Chen et al., 2020) performed wireframe-based GUI retrieval by encoding the visual semantics of UI designs using a large database of UI design wireframes.
@@
By screenshot image: Using a screenshot from an app, it is possible to retrieve similar screenshots from a GUI repository. Screen2vec (Li et al., 2021) extracts the UI components, layout, and app description to generate a screen embedding. Querying similar screenshots is then done by comparing the similarity of screen embeddings. VINS (Bunian et al., 2021) creates screen embeddings by combining both image embeddings and UI components embeddings. Instead of querying screenshots, GUIFetch (Behrang et al., 2018) searches for apps in public repositories that include similar screens and transitions between them.
@@
By text: Using textual queries for GUI retrieval is practical as often only a textual description or keywords are available at the search time. GUIGLE (Bernal-Cardenas et al., 2019) has indexed a collection of GUIs along with their metadata from the ReDraw dataset (Moran et al., 2020). GUIGLE leverages various information sources such as text displayed on the screen, names of UI components, and app name to retrieve relevant screens. On the other hand, as presented in Section 4.2, RaWi (Kolthoff et al., 2023) uses a BERT-based (Devlin et al., 2019) ranking model to retrieve GUIs from the Rico dataset (Deka et al., 2017). This retrieval process involves matching the GUI text (text displayed on the screen, GUI activity name, and GUI component identifier) with the textual query in order to identify the most relevant images. Unlike the preceding studies that focus on retrieving entire screenshot, Gallery D.C. (Chen et al., 2019b; Feng et al., 2022) allows users to search UI components, like a button or a checkbox. It provides a GUI gallery that allows users to search UI components with different filters, including component type, size, colour, app category, and displayed text.
All GUI retrieval approaches suggested so far either use the visual information or the textual information available about the the mobile GUIs for matching the query with corresponding data in the repository. In contract, our approach combines both modalities by training a text-vision model based on screen-caption pairs. This bridges between the visual perception and the natural language and boosts the retrieval performance as our results show.